OpenStudy (anonymous):

complete the patern:6,13,20,27,34 who can complete it???

OpenStudy (amistre64):

you can, i have faith in you :)

OpenStudy (anonymous):

Haha.

OpenStudy (amistre64):

one trick may be to subtract out 6 and see if there is any pattern

OpenStudy (amistre64):

6, 13,20,27,34 -6 -6 -6 -6 -6 -------------- 0, 7, 14, 21, 28 ; these are multiples of 8 right?

OpenStudy (amistre64):

made me typo ... multiples of 7 right? :)

OpenStudy (amistre64):

6 + 7(0) = 6 6 + 7(1) = 13 6 + 7(2) = 20 ....

OpenStudy (anonymous):

grr lol im glad to use this place i do not like math

OpenStudy (amistre64):

An = 6 + 7(n-1) ; if we start at A1 that is

OpenStudy (amistre64):

n: 1 2 3 4 5 6 7 8 A: 6 13 20 27 34 ? ? ? A(n) = 6 + 7(n-1) is a good set up i believe